2. A manufacturer of isopropyl alcohol claims that their product has a mean content of 480 mL. and a standard deviation of 21.5 mL. Assume that the variable is normally distributed. a. If a sample is selected, what is the probability that the content will be less than 505 mL? b. If a sample of 6 isopropyl alcohol is selected, what is the probability that the mean of the sample will be less than 505?


1
Expert's answer
2022-05-03T09:25:08-0400

a.

"P(X<505)=(Z<\frac{505-480}{21.5})"




"=P(Z<1.1628)=0.8775"

b. 

"P(X<505)=(Z<\frac{505-480}{21.5/\sqrt{6}})"




"=P(Z<2.8482)=0.9978"